REU: SCHEDULE FOR WEEKS SEVEN AND EIGHT

MOST TALKS ARE IN ECKHART 206 or 202 (Some may be in R251) (May’s morning talks in R276, noon lunch talks in the barn)

Other early talks for ex–YSP’s and apprentices may be arranged. Ex–YSP’s are expected to arrange morning mentoring sessions with grad students.

Date  SESAME  May  Babai  Hruska  Lewicka  May
Mon, August 1 9–12, 1–4 12:05 – 12:55  4:15 – 6:00
Tue, August 2 9–12, 1–4 12:05 – 12:55 4:15 – 6:00  10:45-11:45
Wed, August 3 9–12, 1–4 12:05 – 12:55  2:00 – 3:30* 4:15 – 6:00 10:45-11:45
Thu, August 4 9–12, 1–4 12:05 – 12:55 4:15 – 6:00  10:45-11:45
Fri, August 5 9–12, 1–4 12:05 – 12:55 2:00 – 3:30* 4:15 – 6:00 10:45-11:45
Mon, August 8 9–12, 1–4 12:05 – 12:55  2:00 – 3:30* 4:15 – 6:00 10:45-11:45
Tue, August 9 9–12, 1–4 12:05 – 12:55 4:15 – 6:00  10:45-11:45
Wed, August 10 9–12, 1–4 12:05 – 12:55  2:00 – 3:30* 4:15 – 6:00 10:45-11:45
Thu, August 11 9–12, 1–4 12:05 – 12:55 4:15 – 6:00  10:45-11:45
Fri, August 12 9–12, 1–4 12:05 – 12:55  2:00 – 3:30 4:15 – 6:00 10:45-11:45

* If there is demand from SESAME participants, Hruska may try to reschedule.

Here is the schedule in summary form:

SEVENTH WEEK:

SESAME 9:00 – 12:00 and 1:00 – 4:00 M-F May (TQFT) 12:05 – 12:50 T-F (bring your own lunch) May (Categories) 10:45 – 11:45 Babai (Potpourri) 4:15 – 6:00 TTh Hruska (Mapping class groups) 4:15 – 6:00 M and 2:00 – 3:30 WF Lewicka (Optimal transportation) 4:15 – 6:00 WF

EIGHTH WEEK:

SESAME 9:00 – 12:00 and 1:00 – 4:00 M-F May (A piacere) 12:05 – 12:50 M-F (bring your own lunch) Babai (Potpourri) 4:15 – 6:00 TTh Hruska (Mapping class groups) 2:00 – 3:30 MWF Lewicka (Optimal transportation) 4:15 – 6:00 MWF

THURSDAY, 6:00 pm: Student presentations and dinner for all!!

NOTE THE CHANGE: Thursday, not Friday, and presentations then.